# End-to-End Probabilistic Framework for Learning with Hard Constraints

### v1

@@ Abstract
We present a general purpose probabilistic forecasting framework, ProbHardE2E, to learn systems that can incorporate operational/physical constraints as hard requirements. ProbHardE2E enforces hard constraints by exploiting variance information in a novel way; and thus it is also capable of performing uncertainty quantification (UQ) on the model. Our methodology uses a novel differentiable probabilistic projection layer (DPPL) that can be combined with a wide range of neural network architectures. This DPPL allows the model to learn the system in an end-to-end manner, compared to other approaches where the constraints are satisfied either through a post-processing step or at inference. In addition, ProbHardE2E can optimize a strictly proper scoring rule, without making any distributional assumptions on the target, which enables it to obtain robust distributional estimates (in contrast to existing approaches that generally optimize likelihood-based objectives, which are heavily biased by their distributional assumptions and model choices); and it can incorporate a range of non-linear constraints (increasing the power of modeling and flexibility). We apply ProbHardE2E to problems in learning partial differential equations with uncertainty estimates and to probabilistic time-series forecasting, showcasing it as a broadly applicable general setup that connects these seemingly disparate domains.
@@ 1 Introduction
The aforementioned hard-constrained models typically generate point estimates, which makes it challenging for them to handle uncertainty. This is particularly problematic in operational or physical domains where uncertainty quantification (UQ) is critical and/or where probabilistic forecasts are needed. Generating output distribution statistics under hard constraints is often computationally expensive or yields only approximate solutions (Robert et al., 1999; Szechtman and Glynn, 2001; Girolami and Calderhead, 2011). There have been domain-specific works in hierarchical probabilistic time series forecasting, which enforce coherency constraints using end-to-end deep learning models (Olivares et al., 2024b; Rangapuram et al., 2021). However, these works either apply only to linear constraints, or they require a computationally expensive sampling procedure in training. Similar approaches have been proposed for computing probabilistic solutions to partial differential equations (PDEs) that satisfy constraints (Hansen et al., 2023; Mouli et al., 2024; Cheng et al., 2025; Gao et al., 2023). In these works, however, the constraints are only applied as a post-processing step, and they do not lead to an end-to-end solution (of interest in the common situation that one wants to incorporate a hard-constrained model within a larger model, and then optimize the larger model) that optimizes the evaluation accuracy. In both the forecasting and PDE application domains, none of this prior UQ work can handle complex nonlinear (hard) constraints.
In this work, we propose a novel probabilistic framework, ProbHardE2E, that can integrate a broad class of hard constraints in an end-to-end fashion, in a way that gracefully incorporates UQ and nonlinear constraints. By leveraging key results from statistics and optimization in a novel way, we predict both the mean and covariance of the output data, moving beyond point estimate predictions. ProbHardE2E enforces nonlinear constraints with an efficient sampling-free method to generate distribution statistics. Among other things, our probabilistic approach enables the effective handling of exogenous spikes and jumps (or other discontinuities) by leveraging data heteroscedasticity, which enhances the model’s robustness and flexibility under varying data conditions.
@@
  • We introduce the ProbHardE2E framework to learn a function in an end-to-end manner by optimizing an objective under hard constraints. The framework enables UQ by learning parameters of a multivariate probabilistic distribution. We show that ProbHardE2E is a general framework that can incorporate a broad class of deep learning backbone models.
@@
[…]
  • We show that ProbHardE2E is effective in two (seemingly-unrelated, but technically-related) tasks, where hard constraints are important: probabilistic time series forecasting; and solving challenging PDEs in scientific machine learning (SciML). We provide an extensive empirical analysis demonstrating that ProbHardE2E results in up to 15×15$\times$15 × lower mean-squared error (MSE) in mean forecast and 2.5×2.5$\times$2.5 × improved variance estimates, measured by the Continuous Ranked Probability Score (CRPS), compared to the baseline methods.
@@ 3 ProbHardE2E: A Unified Probabilistic Optimization Framework
Algorithm 1 ProbHardE2E
0:  Training data $(\phi^{(i)},u^{(i)})\sim\mathcal{D}$, and constraints $g(\cdot)\leq 0$, $h(\cdot)=0$.
0:  Learnable function $\hat{f}_{\theta}:\Phi\to\mathcal{Y}$, which outputs the distribution parameters for $u^{*}\sim\mathbf{Y}_{\theta}$.
@@
3:     Predict unconstrained random variable $\mathbf{Z}_{\theta}(\phi^{(i)})$ in Equation 5. (Predictor Step)
4:     Project $\mathbf{Y}_{\theta}(\phi^{(i)})$ = $\text{DPPL}(\mathbf{Z}_{\theta}(\phi^{(i)}),g(\cdot),h(\cdot))$ in Equation 6. (Corrector Step)
5:     Update $\theta\in\bar{\Theta}$ by minimizing the objective in Equation 4 with the CRPS loss $\ell(\mathbf{Y}_{\theta}({\phi}^{(i)}),u^{(i)})$.
@@ 3.1 Probabilistic Evaluation Metrics and Objective Function
The CRPS naturally aligns with the goal of producing feasible and well-calibrated predictions, as the CRPS rewards distributions that closely match observed outcomes. Enforcing our constraints in the distribution space guarantees that every sample from the predicted distribution is physically or operationally valid. Consequently, modeling the loss through the CRPS provides a principled way to reconcile domain constraints with distributional accuracy.
@@ 3.2 Differentiable Probabilistic Projection Layer (DPPL)
We propose to transform the constrained Problem 1 into the following unconstrained problem:
@@
where $\bar{\Theta}\subseteq\Theta$ denotes the feasible parameter space that ensures constraint satisfaction, and $\ell$ denotes the loss function in Equation 3, through a two-step procedure that: (1) defines a predictive distribution over the outputs; and (2) projects that distribution onto the constraint manifold using a differentiable probabilistic projection layer (DPPL), which guarantees exact constraint satisfaction, in a way that can be optimized end-to-end.
@@
We use an established probabilistic backbone model as the first module of our framework: examples of models that can be used at this step include Gaussian Process (Rasmussen and Williams, 2006), neural process (Kim et al., 2019), DeepVAR (Salinas et al., 2019; Rangapuram et al., 2021), or ensembles of neural networks or operators (Mouli et al., 2024). Let $f_{\theta}:\Phi\rightarrow\mathbb{R}^{k}$ denote this base probabilistic model that predicts the distribution parameters, i.e., the mean $\mu_{\theta}(\phi^{(i)})$ and covariance $\Sigma_{\theta}(\phi^{(i)})$, for $\theta\in\Theta$, but this module is unaware of the constraints. We then use a reparameterization function $r:\mathbb{R}^{k}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ to define the distribution in one of two ways: either as an identity map, where $l=k$, that returns $f_{\theta}(\phi^{(i)})=\big{(}\mu_{\theta}(\phi^{(i)}),\Sigma_{\theta}(\phi^{(i)})\big{)}$ in the case of moment parameterized models (to allow our efficient sample-free paradigm during the training; see Section 3.3); or as a map, where $l=n$, that combines the distribution parameters with noise $\xi\sim p(\xi)$, where $p$ denotes a tractable sampling distribution, and gives a sample from the predicted distribution (to generate constrained samples at inference). This reparameterization function induces the base (unconstrained) distribution parameters or predictive random variable as:
@@
$r(f_{\theta}(\phi^{(i)}),\xi)=\begin{cases}\big{(}\mu_{\theta}(\phi^{(i)}),\Sigma_{\theta}(\phi^{(i)})\big{)},\text{ if $l=k$},\\ \mathbf{Z}_{\theta}(\phi^{(i)}),\text{ if $l=n$}.\end{cases}$ (5)
Observe that our approach is similar to predictor-corrector methods (Boyd and Vandenberghe, 2004; Bertsekas, 1997); and thus we denote this first step as the Predictor Step.
@@
Followed by the backbone model, we propose to use the DPPL to restrict the parameter space to $\bar{\Theta}\subseteq\Theta$, such that for all $\hat{u}_{\theta}(\phi^{(i)})\sim\mathbf{Y}_{\theta}(\phi^{(i)})$, the constraints $g(\hat{u}_{\theta}(\phi^{(i)}))\leq 0$ and $h(\hat{u}_{\theta}(\phi^{(i)}))=0$ are satisfied. The DPPL is our core architecture innovation for leveraging the base model to learn predictions that satisfy the given constraints. We define the projected distribution parameters or projected predictive random variable as:
@@
$\text{DPPL}(r(f_{\theta}(\phi^{(i)}),\xi),g(\cdot),h(\cdot))=r(\hat{f}_{\theta}(\phi^{(i)}),\xi)=\begin{cases}\big{(}\hat{\mu}_{\theta}(\phi^{(i)}),\hat{\Sigma}_{\theta}(\phi^{(i)})\big{)},\text{ if $l=k$},\\ \mathbf{Y}_{\theta}(\phi^{(i)}),\text{ if $l=n$},\end{cases}$ (6)
@@
for $r(f_{\theta}(\phi^{(i)}),\xi)$ in Equation 5, where $\hat{f}_{\theta}:\Phi\rightarrow\mathcal{Y}\subset\mathbb{R}^{k}$ denotes the probabilistic model that outputs the constrained distribution parameters $(\hat{\mu}_{\theta}(\phi^{(i)}),\hat{\Sigma}_{\theta}(\phi^{(i)}))$. Our DPPL yields a constraint-satisfying realization $u^{*}\sim\mathbf{Y}_{\theta}(\phi^{(i)})$ as the final predictive random variable. We denote this second step as the Corrector Step.
@@
Equivalently, the DPPL can be formulated as a constrained least squares problem on the samples of $\mathbf{Z_{\theta}}(\phi^{(i)})$. (See Appendix C for details.) Prior works on imposing hard constraints in time series and solving PDEs (Rangapuram et al., 2021; Hansen et al., 2023) reduce to special cases of our method with linear constraints. (See Appendix D for details.) We draw $z_{\theta}(\phi^{(i)})\sim\mathbf{Z}_{\theta}(\phi^{(i)})$, and we solve the following constrained optimization problem:
@@
$u^{*}(z_{\theta}(\phi^{(i)})):=\operatorname*{argmin}_{\begin{subarray}{c}\hat{u}_{\theta}(\phi^{(i)})\in\mathbb{R}^{n},\\ g(\hat{u}_{\theta}(\phi^{(i)}))\leq 0,\\ h(\hat{u}_{\theta}(\phi^{(i)}))=0\end{subarray}}\lVert\hat{u}_{\theta}(\phi^{(i)})-z_{\theta}(\phi^{(i)})\rVert^{2}_{Q},$ (7)
@@ 3.3 DPPL on the Distribution Parameters for Location-Scale Distributions
In this subsection, we detail how to directly compute the parameters for the constrained distribution by applying our DPPL on the base distribution parameters for a sampling-free approach. To do so, we can assume that the prior distribution $\mathcal{F}$ belongs to a multivariate, location-scale family, i.e., a distribution such that any affine transformation $\mathbf{Y}$ of a random variable $\mathbf{Z}=\mu+\Sigma^{1/2}\xi\sim\mathcal{F}(\mu,\Sigma)$ and $\xi\sim\mathcal{F}(0,1)$, remains within the same distribution family $\mathcal{F}$. Note that this is an example of how to compute the random variable in Equation 5 for a multivariate location-scale distribution. A familiar case of this is when $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ and $\mathbf{Y}=A\mathbf{Z}+B$ is an affine transformation; in which case $\mathbf{Y}\sim\mathcal{N}(A\mu+B,A\Sigma A^{\top})$. Alternatively, we can show that when $\mathbf{Y}$ is a nonlinear transformation of $\mathbf{Z}$, it has approximately (to first-order) the same distribution $\mathbf{Z}$, with an appropriately-chosen set of parameters (given in Equation 8 below). We state this result more formally as the following Theorem 3.1. The proof, given in Appendix F, uses a first-order Taylor expansion to linearize the nonlinear function transformation, and is similar to the Multivariate Delta Method (Casella and Berger, 2001).
@@
Let $\mathbf{Z}\sim\mathcal{F}(\mu,\Sigma)$ be a random variable, where the underlying distribution $\mathcal{F}$ belongs to a multivariate location-scale family of distributions, with mean $\mu$ and covariance $\Sigma$; and let $\mathcal{T}$ be a function with continuous first derivatives, such that ${J_{\mathcal{T}}(\mu)}\Sigma J_{\mathcal{T}}(\mu)^{\top}$ is symmetric positive semi-definite. Then, the transformed distribution $\mathbf{Y}=\mathcal{T}(\mathbf{Z})$ converges in distribution with first-order accuracy to $\mathcal{F}(\hat{\mu},\hat{\Sigma})$ with mean $\hat{\mu}=\mathcal{T}(\mu)$ and covariance $\hat{\Sigma}=J_{\mathcal{T}}(\mu)\Sigma J_{\mathcal{T}}(\mu)^{\top}$, where $J_{\mathcal{T}}(\mu)={\nabla\mathcal{T}(\mu)}^{\top}$ denotes the Jacobian of $\mathcal{T}$ with respect to $z$ evaluated at $\mu$.
Let $\mathbf{Z}\sim\mathcal{F}(\mu,\Sigma)$ denote the prior distribution and $z\sim\mathbf{Z}$. We apply Theorem 3.1 with $\mathcal{T}(z)=u^{*}(z)$, where $u^{*}(z)$ denotes the solution of the constrained least squares problem in Problem 7. In this case, the projected random variable satisfies $\mathbf{Y}\sim\mathcal{F}(\hat{\mu},\hat{\Sigma})$ with updated parameters:
$\hat{\mu}=\mathcal{T}(\mu),$ (8a)
$\hat{\Sigma}=J_{\mathcal{T}}(\mu)\Sigma J_{\mathcal{T}}(\mu)^{\top}.$ (8b)
For practical computation, see Section 3.4, which details how to evaluate $\mathcal{T}(\mu)$ and its Jacobian for different types of constraints.

### v2

@@ Abstract
We present ProbHardE2E, a probabilistic forecasting framework that incorporates hard operational/physical constraints, and provides uncertainty quantification. Our methodology uses a novel differentiable probabilistic projection layer (DPPL) that can be combined with a wide range of neural network architectures. DPPL allows the model to learn the system in an end-to-end manner, compared to other approaches where constraints are satisfied either through a post-processing step or at inference. ProbHardE2E optimizes a strictly proper scoring rule, without making any distributional assumptions on the target, which enables it to obtain robust distributional estimates (in contrast to existing approaches that generally optimize likelihood-based objectives, which are heavily biased by their distributional assumptions and model choices); and it can incorporate a range of non-linear constraints (increasing the power of modeling and flexibility). We apply ProbHardE2E in learning partial differential equations with uncertainty estimates and to probabilistic time-series forecasting, showcasing it as a broadly applicable general framework that connects these seemingly disparate domains.
@@ 1 Introduction
The aforementioned hard-constrained models typically provide point estimates without uncertainty quantification (UQ), limiting their use cases in operational and physical domains requiring probabilistic forecasts. Generating output distribution statistics under hard constraints is often computationally expensive or yields only approximate solutions (Robert et al., 1999; Szechtman and Glynn, 2001; Girolami and Calderhead, 2011). There have been domain-specific works in hierarchical probabilistic time series forecasting, which enforce coherency constraints using end-to-end deep learning models (Olivares et al., 2024b; Rangapuram et al., 2021). However, these works either apply only to linear constraints, or they require a computationally expensive sampling procedure in training. Similar approaches have been proposed for computing probabilistic solutions to partial differential equations (PDEs) that satisfy constraints (Hansen et al., 2023; Mouli et al., 2024; Cheng et al., 2025; Gao et al., 2023; Utkarsh et al., 2025). In these works, however, the constraints are only applied as a post-processing step, and they do not lead to an end-to-end solution (of interest in the common situation that one wants to incorporate a hard-constrained model within a larger model, and then optimize the larger model) that optimizes the evaluation accuracy. In both the forecasting and PDE application domains, none of this prior UQ work can handle complex nonlinear (hard) constraints.
In this work, we propose a novel probabilistic framework, ProbHardE2E, that integrates a broad class of hard constraints (including non-linear constraints) in an end-to-end fashion, while incorporating UQ. By leveraging key results from statistics and optimization in a novel way, we predict both the mean and covariance of the output data, moving beyond point estimate predictions. ProbHardE2E enforces nonlinear constraints with an efficient sampling-free method to generate distribution statistics. Our probabilistic approach enables the effective handling of exogenous spikes and jumps (or other discontinuities) by leveraging data heteroscedasticity, enhancing the model’s robustness and flexibility under varying data conditions.
@@
  • We introduce ProbHardE2E, as a general framework to learn a function in an end-to-end manner by optimizing an objective under hard constraints. The framework enables UQ by learning parameters of a multivariate probabilistic distribution. We show that ProbHardE2E can incorporate a broad class of deep learning backbone models.
@@
  • We use the DPPL to impose constraints directly on the marginals of the multivariate distribution for an efficient sampling-free approach for posterior distribution estimation, which reduces the computational overhead by up to $3\text{--}5\times$ during training.
  • We show that ProbHardE2E is effective in two (seemingly-unrelated, but technically-related) tasks, where hard constraints are important: probabilistic time series forecasting; and solving challenging PDEs in scientific machine learning (SciML). We provide an extensive empirical analysis demonstrating that ProbHardE2E results in up to $15\times$ lower mean-squared error (MSE) in mean forecast and $2.5\times$ improved uncertainty estimates, measured by the Continuous Ranked Probability Score (CRPS), compared to the baseline methods.
@@ 3 ProbHardE2E: A Unified Probabilistic Optimization Framework
Algorithm 1 ProbHardE2E: Training and Inference
0: Training data $\{(\phi^{(i)},u^{(i)})\}\sim\mathcal{D}$, test data $\phi$ and constraints $g(\cdot)\leq 0$, $h(\cdot)=0$.
0: Learnable function $\hat{f}_{\theta}:\Phi\to\mathcal{Y}$ that outputs constrained distribution samples.
@@
3:  Predict unconstrained distribution parameters $(\mu_{\theta}(\phi^{(i)}),\Sigma_{\theta}(\phi^{(i)})).$
4:  Training Mode: Project parameters $(\hat{\mu}_{\theta}(\phi^{(i)}),\hat{\Sigma}_{\theta}(\phi^{(i)}))=\text{DPPL}((\mu_{\theta}(\cdot),\Sigma_{\theta}(\cdot)),g(\cdot),h(\cdot))$.
5:  Update $\theta\in\bar{\Theta}$ by minimizing the CRPS loss $\ell(\mathbf{Y}_{\theta}({\phi}^{(i)}),u^{(i)})$.
@@
7:Inference Mode: Project sample $\mathbf{Y}_{\theta}(\phi)=\text{DPPL}(\mathbf{Z}_{\theta}(\phi),g(\cdot),h(\cdot))$, where $z_{\theta}(\phi)\sim\mathbf{Z}_{\theta}(\phi)$.
8:Return Feasible predicted sample $u^{*}(z_{\theta}(\phi))\sim\mathbf{Y}_{\theta}(\phi)$.
@@ 3.1 Probabilistic Evaluation Metrics and Objective Function
The CRPS naturally aligns with the goal of producing feasible and well-calibrated predictions, as the metric rewards distributions that closely match observed outcomes. Enforcing our constraints in the distribution space guarantees that every sample from the predicted distribution is physically or operationally valid. Consequently, modeling the loss through the CRPS provides a principled way to reconcile domain constraints with distributional accuracy.
@@ 3.2 Differentiable Probabilistic Projection Layer (DPPL)
We transform the constrained Problem 1 into the unconstrained optimization problem:
@@
where $\bar{\Theta}\subseteq\Theta$ denotes the feasible parameter space that ensures constraint satisfaction, and $\ell$ denotes the loss function in Equation 3. We solve this using a two-step procedure: first define a predictive output distribution, then project it onto the constraint manifold using a differentiable probabilistic projection layer (DPPL) for end-to-end optimization.
@@
Our framework begins with an established probabilistic backbone model. This can be a Gaussian Process (Rasmussen and Williams, 2006), neural process (Kim et al., 2019), DeepVAR (Salinas et al., 2019; Rangapuram et al., 2021), or ensembles of neural networks or operators (Mouli et al., 2024). The base model $f_{\theta}:\Phi\rightarrow\mathbb{R}^{k}$ predicts the distribution parameters (mean $\mu_{\theta}(\phi^{(i)})$ and covariance $\Sigma_{\theta}(\phi^{(i)})$, for $\theta\in\Theta$) – without constraint awareness. We then use a reparameterization function $r:\mathbb{R}^{k}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ to define the distribution in one of two ways: either as an identity map, where $l=k$, that returns $f_{\theta}(\phi^{(i)})=\big(\mu_{\theta}(\phi^{(i)}),\Sigma_{\theta}(\phi^{(i)})\big)$ for our efficient sample-free paradigm during training; or as a map, where $l=n$, that combines the distribution parameters with noise $\xi\sim p(\xi)$, where $p$ denotes a tractable sampling distribution, and gives a sample $z_{\theta}(\phi^{(i)})\sim\mathbf{Z}_{\theta}(\phi^{(i)})$ from the predicted distribution to generate constrained samples at inference. This dual-mode design balances training efficiency with strict constraint feasibility at inference.
The reparameterization function induces the base (unconstrained) distribution parameters or predictive random variable as:
@@
$r(f_{\theta}(\phi^{(i)}),\xi)=\begin{cases}\big(\mu_{\theta}(\phi^{(i)}),\Sigma_{\theta}(\phi^{(i)})\big),\hfill\text{ (Training)}\\ \mathbf{Z}_{\theta}(\phi^{(i)}),\hfill\text{ (Inference)}\end{cases}$ (5)
@@
Following this Predictor Step above, we use the DPPL in the Corrector Step to restrict the parameter space to $\bar{\Theta}\subseteq\Theta$, such that for all $\hat{u}_{\theta}(\phi^{(i)})\sim\mathbf{Y}_{\theta}(\phi^{(i)})$, the constraints $g(\hat{u}_{\theta}(\phi^{(i)}))\leq 0$ and $h(\hat{u}_{\theta}(\phi^{(i)}))=0$ are satisfied. The DPPL is our core architecture innovation for leveraging the base model to learn predictions that satisfy the given constraints. We define the projected distribution parameters or projected predictive random variable as:
@@
$\text{DPPL}(r(f_{\theta}(\phi^{(i)}),\xi),g(\cdot),h(\cdot))=r(\hat{f}_{\theta}(\phi^{(i)}),\xi)=\begin{cases}\big(\hat{\mu}_{\theta}(\phi^{(i)}),\hat{\Sigma}_{\theta}(\phi^{(i)})\big),\hfill\text{ (Training)}\\ \mathbf{Y}_{\theta}(\phi^{(i)}),\hfill\text{ (Inference)}\end{cases}$ (6)
@@
for $r(f_{\theta}(\phi^{(i)}),\xi)$ in Equation 5, where $\hat{f}_{\theta}:\Phi\rightarrow\mathcal{Y}\subset\mathbb{R}^{k}$ denotes the probabilistic model that outputs the constrained distribution parameters $(\hat{\mu}_{\theta}(\phi^{(i)}),\hat{\Sigma}_{\theta}(\phi^{(i)}))$. Our DPPL yields a constraint-satisfying realization $u^{*}\sim\mathbf{Y}_{\theta}(\phi^{(i)})$ as the final predictive random variable.
@@
This two-step approach mirrors predictor-corrector methods (Boyd and Vandenberghe, 2004; Bertsekas, 1997), with the DPPL serving as our key architectural innovation for ensuring constraint satisfaction. Equivalently, the DPPL can be formulated as a constrained least squares problem on the samples of $\mathbf{Z_{\theta}}(\phi^{(i)})$. (See Appendix C for details.) Prior works on imposing hard constraints in time series and solving PDEs (Rangapuram et al., 2021; Hansen et al., 2023) reduce to special cases of our method with linear constraints. (See Appendix D for details.) We draw $z_{\theta}(\phi^{(i)})\sim\mathbf{Z}_{\theta}(\phi^{(i)})$, and we solve the following constrained optimization problem:
@@
$u^{*}(z_{\theta}(\phi^{(i)})):=\operatorname*{argmin}_{\begin{subarray}{c}\hat{u}_{\theta}(\phi^{(i)})\in\mathbb{R}^{n},g(\hat{u}_{\theta}(\phi^{(i)}))\leq 0,h(\hat{u}_{\theta}(\phi^{(i)}))=0\end{subarray}}\lVert\hat{u}_{\theta}(\phi^{(i)})-z_{\theta}(\phi^{(i)})\rVert^{2}_{Q},$ (7)
@@ 3.3 DPPL on the Distribution Parameters for Location-Scale Distributions
In this subsection, we detail how to directly compute the parameters for the constrained distribution by applying our DPPL on the base distribution parameters for an efficient, sampling-free during training. To do so, we can assume that the prior distribution $\mathcal{F}$ belongs to a multivariate, location-scale family, i.e., a distribution such that any affine transformation $\mathbf{Y}$ of a random variable $\mathbf{Z}=\mu+\Sigma^{1/2}\xi\sim\mathcal{F}(\mu,\Sigma)$ and $\xi\sim\mathcal{F}(0,1)$, remains within the same distribution family $\mathcal{F}$. This is an example of how to compute the random variable in Equation 5 for a multivariate location-scale distribution. A familiar case of this is when $\mathbf{Z}\sim\mathcal{N}(\mu,\Sigma)$ and $\mathbf{Y}=A\mathbf{Z}+B$ is an affine transformation; in which case $\mathbf{Y}\sim\mathcal{N}(A\mu+B,A\Sigma A^{\top})$. Alternatively, we can show that when $\mathbf{Y}$ is a nonlinear transformation of $\mathbf{Z}$, it has approximately (to first-order) the same distribution $\mathbf{Z}$, with an appropriately-chosen set of parameters (given in Equation 8 below). We state this result more formally in Theorem 3.1. The proof, given in Appendix F, uses a first-order Taylor expansion to linearize the nonlinear function transformation, and is similar to the Multivariate Delta Method (Casella and Berger, 2001).
@@
[…]
$\hat{\mu}=\mathcal{T}(\mu),\qquad\hat{\Sigma}=J_{\mathcal{T}}(\mu)\,\Sigma\,J_{\mathcal{T}}(\mu)^{\top}.$ (8)